# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v1

@@ 5 Sequential convergence under KL property
$\partial E_{p}(z,u)=\begin{pmatrix}\nabla_{x}G_{p}(z)-u+\mathcal{N}_{X}(x)\\ \partial_{y}G_{p}(z)+\nabla_{y}H(u,y)+\mathcal{N}_{Y}(y)\\ \nabla_{u}H(u,y)-x\end{pmatrix}$ (60)
Using the update rules of $x^{k+1}$ and $y^{k+1}$ in (20) and (15), along with the formulas for $d_{x}^{k}$ and $d^{k}_{y}$ in (19) and (14), we have
@@
00 0 $\in\frac{1}{p_{k}}\nabla_{y}F(z^{k})+\nabla_{y}f(z^{k})+\partial_{y}g(x^{k},y^{k+1})-\frac{y^{k}-\theta^{k}}{\gamma}+\frac{y^{k+1}-y^{k}}{\beta_{k}}+\mathcal{N}_{Y}(y^{k+1}).$
@@
$-\nabla_{y}v_{\gamma}(x^{k},y^{k+1})+(y^{k}-\theta^{k})/\gamma-(y^{k+1}-y^{k})/\beta_{k}.$
@@
$\leq$ $\left(L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+1/\underline{\alpha}+\rho_{v_{1}}\right)\|x^{k+1}-x^{k}\|+(L_{f_{x}}+L_{g_{2}})C_{\eta}s_{k}$
@@
$\leq$ $\left(L_{F_{y}}/p_{k}+L_{f_{y}}+L_{g_{y}}+L_{\theta^{*}}/\gamma+1/\underline{\beta}+1/\gamma\right)\|z^{k+1}-z^{k}\|+{C_{\eta}}s_{k}/\gamma,$
where the last inequality follows from (30) and the $L_{\theta^{*}}$-Lipschitz continuity of $\theta^{*}(x,y)$ with respect to $y$ (from Lemma 2). For the last component of $\partial E_{p_{k}}(z^{k+1},u^{k})$, note that since $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$, we have $\nabla_{u}H(u^{k},y^{k+1})=x^{k}$. Thus
@@
for some $p_{s}<2$. For example, $s_{k}$ can be chosen as $s_{k}=1/k^{1.1}$. Under this condition, there exists $q>0$ such that $2(1-1/q)>p_{s}$, ensuring that $\sum_{k=1}^{\infty}ks_{k}^{2(1-1/q)}<\infty$. To facilitate the convergence analysis, we introduce the following auxiliary function,
@@
where $E_{p}(z,u)$ is as defined in (51). Since $w^{q}$ is a KL function, it follows from [56, Theorem 3.6] (although this result considers the generalized concave KL property, it can be directly extended to the KL property using the same proof) that if $E_{p}(z,u)$ is a KL function, then $\tilde{E}_{p}(z,u,w)$ is also a KL function.
@@
Under Assumptions 1, 2, 3 and 4, let $\epsilon>0$. Consider the sequences $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ generated by AGILS (1) under the same conditions as in Section 4.2, with $s_{k}$ satisfying condition (66). Suppose the sequence $\left\{(x^{k},y^{k})\right\}$ is bounded and $p_{k}=\bar{p}$ for sufficiently large $k$ and $E_{\bar{p}}(z,u)$ is a KL function. Then the sequence $\left\{x^{k},y^{k}\right\}$ converges to a KKT point of $({\rm VP})_{\gamma}^{\tilde{\epsilon}}$ for some $\tilde{\epsilon}\leq\epsilon$.
@@
Since there exists $k_{0}>0$ such that $p_{k}=\bar{p}:=p_{k_{0}}$ for any $k\geq k_{0}$. Without loss of generality, we assume that $p_{k}=\bar{p}$ for all $k$. From Lemma 6, we know that $\lim_{k\rightarrow\infty}\|z^{k+1}-z^{k}\|^{2}=0$. Let $\Omega$ denote the set of all limit points of the sequence $\{(z^{k},u^{k-1})\}$. From Proposition 2, we know that $\nabla v_{\gamma}$ is continuous. Given that the sequence $\{z^{k}\}$ is bounded, and $u^{k}=\nabla_{x}v_{\gamma}(x^{k},y^{k+1})+\rho_{v_{1}}x^{k}$, we conclude that the sequence $\{u^{k}\}$ is also bounded. Consequently, $\Omega$ is a bounded set. Since limit points form a closed set, it follows that $\Omega$ is compact and $\lim_{k\rightarrow\infty}\mathrm{dist}((z^{k},u^{k-1}),\Omega)=0$. Now, define $w^{k}=(\sum_{i=k}^{\infty}\nu_{i})^{1/q}$, since $\sum_{k}\nu_{k}<\infty$, we have $w^{k}\rightarrow 0$. We can also have $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})=E_{\bar{p}}(z^{k},u^{k-1})+\sum_{i=k}^{\infty}\nu_{i}$. By Lemma 8, there exist constants $a,b>0$ such that
@@
${\rm dist}\left(0,\tilde{E}_{\bar{p}}\left(z^{k+1},u^{k},w^{k+1}\right)\right)\leq b\|z^{k+1}-z^{k}\|+\tilde{\nu}_{k}+q|w^{k+1}|^{q-1}.$ (68)
Since $\tilde{E}_{\bar{p}}(z,u,w)$ is bounded below and continuous on $X\times Y$, $\lim_{k\rightarrow\infty}\|z^{k+1}-z^{k}\|=0$, and $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})$ is nonincreasing as $k$ increases, therefore, there exists a constant $\bar{E}$ such that for any subsequence $\{(z^{k_{j}},u^{k_{j}-1},w^{k_{j}})\}$ of sequence $\{(z^{k},u^{k-1},w^{k})\}$, we have $\bar{E}=\lim_{k_{j}\rightarrow\infty}\tilde{E}_{\bar{p}}(z^{k_{j}},$ $u^{k_{j}-1},w^{k_{j}})=\lim_{k\rightarrow\infty}\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})$. This implies that the function $\tilde{E}_{\bar{p}}(z,u,w)$ is constant on $\Omega\times\{0\}$. We can assume that $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})>\bar{E}$ for all $k$. If there exists some $k>0$ such that $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})=\bar{E}$, then from (67), we would have $\|z^{k+1}-z^{k}\|=0$ for sufficiently large $k$, implying that the sequence $\{z^{k}\}$ converges and the conclusion follows immediately.
@@
$\sum_{k=k_{0}}^{K}\|z^{k+1}-z^{k}\|\leq\|z^{1}-z^{0}\|+\frac{b}{a}\phi\left(\tilde{E}_{1}\right)+\sum_{k=k_{0}}^{K}\frac{\tilde{\nu}_{k-1}}{b}+\sum_{k=k_{0}}^{K}\frac{q|w^{k}|^{q-1}}{b}.$
Taking $K\rightarrow\infty$ in the above inequality, and using the fact that $\sum_{k=1}^{\infty}\tilde{\nu}_{k-1}<\infty$ and $\sum_{k=1}^{\infty}|w^{k}|^{q-1}=\sum_{k=1}^{\infty}|(\sum_{i=k}^{\infty}\nu_{i})^{1/q}|^{q-1}\leq\sum_{k=1}^{\infty}\sum_{i=k}^{\infty}\nu_{i}^{1-1/q}=\sum_{k=1}^{\infty}k\nu_{i}^{1-1/q}\leq C_{\nu}(\sum_{k=1}^{\infty}ks_{k}^{2(1-1/q)})<\infty$ for some $C_{\nu}>0$, we obtain
$\sum_{k=1}^{\infty}\|z^{k+1}-z^{k}\|<\infty.$
Thus, the sequence $\{z^{k}\}$ is a Cauchy sequence. Hence the sequence $\{z^{k}\}$ converges and we get the conclusion from Theorem 2
@@
To further substantiate the applicability of our sequential convergence theory in applications, we now discuss the KL property imposed on the merit function $E_{\bar{p}}(z,u)$ in Theorem 5. A wide range of functions automatically satisfy the KL property, see, e.g., [57, 58, 59, 52]. Notably, as shown in [60, 61], semi-algebraic function is an important subclass of KL functions. More importantly, the class of semi-algebraic functions is closed under various operations, see, e.g., [57, 58, 52]. Specifically, the indicator functions of semi-algebraic sets, finite sum and product of semi-algebraic functions, composition of semi-algebraic functions and partial minimization of semi-algebraic function over semi-algebraic set are all semi-algebraic functions.
@@ 6 Numerical experiments
In this section, we evaluate the performance of our proposed algorithm, AGILS, by comparing it with several existing methods on two problems: a toy example and the sparse group Lasso bilevel hyperparameter selection problem. Both problems are specific instances of bilevel problem (1) and satisfy Assumptions 1, 2, 3, and 4. The baseline methods for comparison include: grid search, random search, Tree-structured Parzen Estimator (TPE [62] [^1] https://github.com/hyperopt/hyperopt, a Bayesian optimization method), Implicit Gradient-based Joint Optimization (IGJO [32] [^2] https://github.com/jjfeng/nonsmooth-joint-opt, an implicit differentiation method), and Moreau Envelope-based Hessian-free Algorithm (MEHA [43], a single-loop, gradient-based method), Value Function-Based Difference-of-Convex Algorithm (VF-iDCA [28] [^3] https://github.com/SUSTech-Optimization/VF-iDCA, a method based on a difference-of-convex algorithm ), and the MPCC approach (which reformulates the bilevel problem as a mathematical program with complementarity constraints). Methods that require smoothness assumptions, such as those in [34, 39], were excluded because the tested problems do not satisfy the required smoothness conditions. For grid search, random search, and TPE, each subproblem was solved using the CVXPY package with the CLARABEL and SCS solvers. All experiments were implemented in Python and conducted on a laptop with an Intel i7-1260P CPU (2.10 GHz) and 32 GB RAM.
@@ 6.1 Toy example
$\min\limits_{x\in{\mathbb{R}}^{n},y\in{\mathbb{R}}^{n},0\leq x\leq 1}\sum\limits_{i=0}^{n}y_{i}\quad{\rm s.t.}~{}~{}y\in\underset{y\in{\mathbb{R}}^{n}}{\mathrm{argmin}}\ \sum\limits_{i=0}^{n}\sqrt{(y_{i}-a_{i})^{2}+\frac{1}{n^{2}}}+\sum\limits_{i=0}^{n}x_{i}|y_{i}|,$ (69)
@@
where $a\in{\mathbb{R}}^{n}$ satisfies $a_{i}=-2/n^{2/3}$ for $i\in 1,\dots,n/2$, and $a_{i}=-2/n^{2/3}$ for $i\in n/2+1,\ldots,n$. The optimal solution is $x^{*}_{i}=0$, $y^{*}_{i}=a_{i}$ for $i=1,\dots,n/2$ and $x^{*}_{i}\in[a_{i}/\sqrt{a_{i}^{2}+1/n^{2}},1]$, $y^{*}_{i}=0$ for $i=n/2+1,\dots,n$.
@@
We compare the performance of AGILS with several methods, including grid search, random search, TPE, MEHA, and the MPCC approach. The performance of each method is evaluated using metric Error $:={\rm dist}(z^{k},S^{*})/\sqrt{1+\min_{z\in S^{*}}\lVert z^{*}\rVert^{2}}$, where $S^{*}$ is the optimal set of (69), and $z:=(x,y)$. For grid search, we set $x_{1}=x_{2}=\dots=x_{n}=\rho$, and perform the search over $\rho$ on a grid of 100 uniformly spaced values within the interval $[0,1]$. Random search is implemented by uniformly sampling 100 points from $[0,1]^{n}$. For TPE, the search space is modeled as a uniform distribution over $[0,1]^{n}$. In the MPCC approach, we first reformulate the toy example by introducing non-negative vectors $u$ and $v$ to represent $y=u-v$, and transform $\|y\|_{1}$ into $\sum_{i=1}^{n}(u_{i}+v_{i})$. Then the corresponding MPCC reformulation is solved using the IPOPT solver with a relaxation parameter of 0.1 on the complementarity constraint and a solver tolerance of 0.1. For MEHA, the initial iterates are the same as AGILS, and the parameters are searched over $\tilde{c}_{0}\in\{0.1,0.5,1,5\}$, $p\in\{0.1,0.2,0.3,0.4\}$ and $\tilde{\alpha}$, $\tilde{\beta}$, $\tilde{\eta}$, and $\tilde{\gamma}$ scaled proportionally from the AGILS parameters ${\alpha}$, ${\beta}$, ${\eta}$, and ${\gamma}$ using scales $\{1,1/2,1/3,1/4\}$. The optimal parameters that minimize the Error are selected as $\tilde{c}_{0}=1$, $p=0.2$, $\tilde{\gamma}=\gamma/3$, $\tilde{\alpha}=\alpha/4$, $\tilde{\beta}=\beta$, and $\tilde{\eta}=\eta$. MEHA terminates when $\mathrm{Error}<1/n$. For AGILS, the initial iterates are $x_{0}=y_{0}=\theta_{0}=[0,\dots,0]$, the parameters are set as $\epsilon=10^{-6}$, $s_{k}=0.05/(k+1)^{1.05},$ and $p_{0}=0.5,\ \varrho_{p}=0.02,\ c_{p}=1$. The step sizes $\alpha,\beta,\gamma$ are chosen as $\alpha_{k}=1/(L_{\psi_{x,k}}+c_{\alpha})$, $\beta_{k}=1/(L_{\psi_{y,k}}+c_{\beta})$, and $\gamma=1/(\rho_{f_{2}}+\rho_{g_{2}})$ where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ with the following values for the constants $L_{F_{x}}=L_{F_{y}}=L_{f_{x}}=\rho_{f_{1}}=\rho_{f_{2}}=L_{g_{1}}=0$, $\rho_{g_{1}}=\rho_{g_{2}}=1$, $L_{f_{y}}=n$, $c_{\alpha}=c_{\beta}=0.1$ and $\eta=1/(L_{f_{y}}+1/\gamma)$. AGILS terminates when $\text{Error}<1/n$. The approximations $\theta^{k}$ are found at each iteration using the proximal gradient method to satisfy inexact criteria (18) and (21).
@@
Numerical results for two different problem dimensions, $n=200$ and $n=600$, are reported in Table 1. The results demonstrate that AGILS performs well in both computational efficiency and solution accuracy, consistently achieving the lowest Error with the shortest computational time. MEHA performs second, but it requires tuning of parameters from a set of candidate values, which can be a limitation.
@@
Table 1: Comparisons of different methods on the toy example for dimensions $n=200$ and $n=600$
@@
Method Grid Random TPE MPCC MEHA AGILS
Time (s) 0.62 4.83 28.55 19.23 0.08 0.08
Error 0.70 0.75 0.72 0.02 0.00 0.00

### v2

@@ 5 Sequential convergence under KL property
$\partial E_{p}(z,u)=\begin{pmatrix}\nabla_{x}G_{p}(z)-u+\mathcal{N}_{X}(x)\\ \partial_{y}G_{p}(z)+\nabla_{y}H(u,y)+\mathcal{N}_{Y}(y)\\ \nabla_{u}H(u,y)-x\end{pmatrix}.$ (60)
Using the update rules of $x^{k+1}$ and $y^{k+1}$ in (19) and (14) together with the assumption $\tilde{y}^{k}=y^{k}$, and the formulas for $d_{x}^{k}$ and $d^{k}_{y}$ in (18) and (13), we have
@@
00 0 $\in\frac{1}{p_{k}}\nabla_{y}F(z^{k})+\nabla_{y}f(z^{k})+\partial_{y}g(x^{k},y^{k+1})-\frac{\tilde{y}^{k}-\theta^{k}}{\gamma}+\frac{y^{k+1}-\tilde{y}^{k}}{\beta_{k}}+\mathcal{N}_{Y}(y^{k+1}).$
@@
$-\nabla_{y}v_{\gamma}(x^{k},y^{k+1})+(\tilde{y}^{k}-\theta^{k})/\gamma-(y^{k+1}-\tilde{y}^{k})/\beta_{k}.$
@@
$\leq$ $\left(L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+1/\underline{\alpha}+\rho_{v_{1}}\right)\|x^{k+1}-x^{k}\|+(L_{f_{x}}+L_{g_{2}})C_{\eta}\zeta_{k}$
@@
$\leq$ $\left(L_{F_{y}}/p_{k}+L_{f_{y}}+L_{g_{y}}+L_{\theta^{*}}/\gamma+1/\underline{\beta}+1/\gamma\right)\|z^{k+1}-z^{k}\|+{C_{\eta}}\zeta_{k}/\gamma,$
where the last inequality follows from (34) and the $L_{\theta^{*}}$-Lipschitz continuity of $\theta^{*}_{\gamma}(x,y)$ with respect to $y$ (from Lemma 1). For the last component of $\partial E_{p_{k}}(z^{k+1},u^{k})$, note that since $u^{k}=\nabla_{x}\tilde{v}_{\gamma}(x^{k},y^{k+1})$, we have $\nabla_{u}H(u^{k},y^{k+1})=x^{k}$. Thus
@@
for some $p_{s}<2$. For example, $s_{k}$ can be chosen as $s_{k}=1/k^{1.1}$. Under these conditions, we can choose a sufficiently large $q>0$ such that $2(1-1/q)>p_{s}$, ensuring that $\sum_{k=1}^{\infty}ks_{k}^{2(1-1/q)}<\infty$. Furthermore, by applying the ratio test, as in the proof of Lemma 5, we can show that $\sum_{k=1}^{\infty}k(\prod_{i=0}^{k}\tau_{i})\mathcal{G}(\theta^{0},x^{0},y^{0})^{2(1-1/q)}<\infty$. Since $\zeta_{k}:=\max\{s_{k},(\prod_{i=0}^{k}\tau_{i})\mathcal{G}(\theta^{0},x^{0},y^{0})\}$, it follows that $\sum_{k=1}^{\infty}k\zeta_{k}^{2(1-1/q)}<\infty$.
The main difficulty in the convergence analysis arises from the inequality (53). The presence of the error term $\nu_{k}$ means this condition does not enforce a sufficient decrease in the merit function $E_{p}$ at each iteration. To address this challenge, we introduce the following auxiliary merit function for $E_{p}(z,u)$,
@@
where $E_{p}(z,u)$ is as defined in (52). Since $w^{q}$ is a KL function, it follows from [58, Theorem 3.6] (although this result considers the generalized concave KL property, it can be directly extended to the KL property using the same proof) that if $E_{p}(z,u)$ is a KL function, then $\tilde{E}_{p}(z,u,w)$ is also a KL function.
@@
Under Assumptions 1, 2, 3 and 4, let $\epsilon>0$. Consider the sequences $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ generated by AGILS (Algorithm 1) under the same conditions as in Section 4.2, with $s_{k}$ satisfying condition (66). Suppose the sequence $\{(x^{k},y^{k})\}$ is bounded, and for sufficiently large $k$, $p_{k}=\bar{p}$ and the feasibility correction procedure (Algorithm 2) is not applied ( i.e., $\tilde{y}^{k}=y^{k}$). If $E_{\bar{p}}(z,u)$ is a KL function, then the sequence $\left\{x^{k},y^{k}\right\}$ converges to a KKT point of $({\rm VP})_{\gamma}^{\tilde{\epsilon}}$ for some $\tilde{\epsilon}\leq\epsilon$.
@@
By assumption, the penalty parameter $p_{k}$ is eventually constant and the feasibility correction procedure is eventually inactive. Therefore, there exists $k_{0}\in\mathbb{N}$ such that for all $k\geq k_{0}$, we have $p_{k}=\bar{p}:=p_{k_{0}}$, and $\tilde{y}^{k}=y^{k}$. From Lemma 6, we know that $\lim_{k\rightarrow\infty}\|z^{k+1}-z^{k}\|^{2}=0$. Let $\Omega$ denote the set of all limit points of the sequence $\{(z^{k},u^{k-1})\}$. From Proposition 2, we know that $\nabla v_{\gamma}$ is continuous. Given that the sequence $\{z^{k}\}$ is bounded, and $u^{k}=\nabla_{x}v_{\gamma}(x^{k},y^{k+1})+\rho_{v_{1}}x^{k}$, we conclude that the sequence $\{u^{k}\}$ is also bounded. Consequently, $\Omega$ is a bounded set. Since limit points form a closed set, it follows that $\Omega$ is compact and $\lim_{k\rightarrow\infty}\mathrm{dist}((z^{k},u^{k-1}),\Omega)=0$. Now, define $w^{k}=(\sum_{i=k}^{\infty}\nu_{i})^{1/q}$, since $\sum_{k}\nu_{k}<\infty$, we have $w^{k}\rightarrow 0$. We can also have $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})=E_{\bar{p}}(z^{k},u^{k-1})+\sum_{i=k}^{\infty}\nu_{i}$. By Lemma 9, there exist constants $a,b>0$ such that
@@
${\rm dist}\left(0,\partial\tilde{E}_{\bar{p}}\left(z^{k+1},u^{k},w^{k+1}\right)\right)\leq b\|z^{k+1}-z^{k}\|+\tilde{\nu}_{k}+q|w^{k+1}|^{q-1}.$ (68)
Since $\tilde{E}_{\bar{p}}(z,u,w)$ is bounded below and continuous on $X\times Y\times{\mathbb{R}}_{+}$, $\lim_{k\rightarrow\infty}\|z^{k+1}-z^{k}\|=0$, $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})$ is nonincreasing as $k$ increases, therefore, there exists a constant $\bar{E}$ such that for any subsequence $\{(z^{k_{j}},u^{k_{j}-1},w^{k_{j}})\}$ of the sequence $\{(z^{k},u^{k-1},w^{k})\}$, we have $\bar{E}=\lim_{k_{j}\rightarrow\infty}\tilde{E}_{\bar{p}}(z^{k_{j}},$
$u^{k_{j}-1},w^{k_{j}})=\lim_{k\rightarrow\infty}\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})$. This implies that the function $\tilde{E}_{\bar{p}}(z,u,w)$ is constant on $\Omega\times\{0\}$. We can assume that $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})>\bar{E}$ for all $k$. If there exists some $k>0$ such that $\tilde{E}_{\bar{p}}(z^{k},u^{k-1},w^{k})=\bar{E}$, then from (67), we would have $\|z^{k+1}-z^{k}\|=0$ for sufficiently large $k$, implying that the sequence $\{z^{k}\}$ converges and the conclusion follows immediately.
@@
$\sum_{k=k_{0}}^{K}\|z^{k+1}-z^{k}\|\leq\|z^{k_{0}}-z^{k_{0}-1}\|+\frac{b}{a}\phi\left(\tilde{E}_{k_{0}}\right)+\sum_{k=k_{0}}^{K}\frac{\tilde{\nu}_{k-1}}{b}+\sum_{k=k_{0}}^{K}\frac{q|w^{k}|^{q-1}}{b}.$
Taking $K\rightarrow\infty$ in the above inequality, and using the fact that $\sum_{k=1}^{\infty}\tilde{\nu}_{k-1}<\infty$ and $\sum_{k=1}^{\infty}|w^{k}|^{q-1}=\sum_{k=1}^{\infty}|(\sum_{i=k}^{\infty}\nu_{i})^{1/q}|^{q-1}\leq\sum_{k=1}^{\infty}\sum_{i=k}^{\infty}\nu_{i}^{1-1/q}=\sum_{k=1}^{\infty}k\nu_{k}^{1-1/q}\leq C_{\nu}(\sum_{k=1}^{\infty}k\zeta_{k}^{2(1-1/q)})<\infty$ for some $C_{\nu}>0$, we obtain $\sum_{k=1}^{\infty}\|z^{k+1}-z^{k}\|<\infty$. Thus, the sequence $\{z^{k}\}$ is a Cauchy sequence. Hence the sequence $\{z^{k}\}$ converges and we get the conclusion from Theorem 2. ∎
@@
To further substantiate the applicability of our sequential convergence theory in applications, we now discuss the KL property imposed on the merit function $E_{\bar{p}}(z,u)$ in Theorem 4. A wide range of functions automatically satisfy the KL property, see, e.g., [3, 4, 13, 14]. Notably, as shown in [11, 12], semi-algebraic function is an important subclass of KL functions. More importantly, the class of semi-algebraic functions is closed under various operations, see, e.g., [3, 4, 14]. Specifically, the indicator functions of semi-algebraic sets, finite sum and product of semi-algebraic functions, composition of semi-algebraic functions and partial minimization of semi-algebraic functions over semi-algebraic sets are all semi-algebraic functions.
@@ 6 Numerical experiments
In this section, we evaluate the performance of our proposed algorithm, AGILS, by comparing it with several existing methods on two problems: a toy example and the sparse group Lasso bilevel hyperparameter selection problem. AGILS utilizes an inexactness criterion that can be either absolute or relative. To explore this further, we also examine two variants: AGILS_A, which exclusively employs an absolute inexactness criterion for approximating $\theta$, and AGILS_R, which uses only a relative inexactness criterion for the same purpose. Both problems are specific instances of bilevel problem (1) and satisfy Assumptions 1, 2, 3, and 4. The baseline methods for comparison include: grid search, random search, Tree-structured Parzen Estimator (TPE [9] [^1] https://github.com/hyperopt/hyperopt, a Bayesian optimization method), Implicit Gradient-based Joint Optimization (IGJO [26] [^2] https://github.com/jjfeng/nonsmooth-joint-opt, an implicit differentiation method), and Moreau Envelope-based Hessian-free Algorithm (MEHA [42], a single-loop, gradient-based method), Value Function-Based Difference-of-Convex Algorithm (VF-iDCA [30] [^3] https://github.com/SUSTech-Optimization/VF-iDCA, a method based on a difference-of-convex algorithm), and the MPCC approach (which reformulates the bilevel problem as a mathematical program with complementarity constraints). Methods that require smoothness assumptions, such as those in [34, 44], were excluded because the tested problems do not satisfy the required smoothness conditions. For grid search, random search, and TPE, each subproblem was solved using the CVXPY package with the CLARABEL and SCS solvers. All experiments were implemented in Python and conducted on a laptop with an Intel i7-1260P CPU (2.10 GHz) and 32 GB RAM.
@@ 6.1 Toy example
$\min\limits_{x\in{\mathbb{R}}^{n},y\in{\mathbb{R}}^{n},0\leq x\leq 1}\sum\limits_{i=1}^{n}y_{i}\quad{\rm s.t.}~{}~{}y\in\underset{y\in{\mathbb{R}}^{n}}{\mathrm{argmin}}\ \sum\limits_{i=1}^{n}\sqrt{(y_{i}-a_{i})^{2}+\frac{1}{n^{2}}}+\sum\limits_{i=1}^{n}x_{i}|y_{i}|,$ (69)
@@
where $a\in{\mathbb{R}}^{n}$ satisfies $a_{i}=-2/n^{2/3}$ for $i\in 1,\dots,n/2$, and $a_{i}=2/n^{2/3}$ for $i\in n/2+1,\ldots,n$. The optimal solution is $x^{*}_{i}=0$, $y^{*}_{i}=a_{i}$ for $i=1,\dots,n/2$ and $x^{*}_{i}\in[a_{i}/\sqrt{a_{i}^{2}+1/n^{2}},1]$, $y^{*}_{i}=0$ for $i=n/2+1,\dots,n$.
@@
We compare the performance of AGILS and its two variants with several methods, including grid search, random search, TPE, IGJO, VF-iDCA, MEHA, and the MPCC approach. The performance of each method is evaluated using metric Error $:={\rm dist}(z^{k},S^{*})/\sqrt{1+\min_{z\in S^{*}}\lVert z^{*}\rVert^{2}}$, where $S^{*}$ is the optimal set of (69), and $z:=(x,y)$. For grid search, we set $x_{1}=x_{2}=\dots=x_{n}=\rho$, and perform the search over $\rho$ on a grid of 100 uniformly spaced values within the interval $[0,1]$. Random search is implemented by uniformly sampling 100 points from $[0,1]^{n}$. For TPE, the search space is modeled as a uniform distribution over $[0,1]^{n}$. In the MPCC approach, we first reformulate the toy example by introducing non-negative vectors $u$ and $v$ to represent $y=u-v$, and transform $\|y\|_{1}$ into $\sum_{i=1}^{n}(u_{i}+v_{i})$. The resulting MPCC is solved using IPOPT, a primal-dual interior point solver for large-scale nonlinear programs, with a relaxation parameter and tolerance of 0.1. All algorithms (VF-iDCA, MEHA, IGJO) use the same initial iterates as AGILS. For VF-iDCA, we set $\beta_{0}=0.5$, $\rho=0.5$, $c=5$, and $\delta=5$. Since VF-iDCA and IGJO may terminate without converging to a true solution, we impose additional stopping rules to ensure numerical reliability. VF-iDCA stops when $\lVert z^{k+1}-z^{k}\rVert/\sqrt{1+\lVert z^{k}\rVert^{2}}<10^{-4}$ or $\mathrm{Error}<1/n$, while IGJO stops when $\lVert x^{k+1}-x^{k}\rVert/\sqrt{1+\lVert x^{k}\rVert^{2}}<10^{-4}$ or until $\mathrm{Error}<1/n$. For MEHA, and the parameters are searched over $\tilde{c}_{0}\in\{0.1,0.5,1,5\}$, $p\in\{0.1,0.2,0.3,0.4\}$ and $\tilde{\alpha}$, $\tilde{\beta}$, $\tilde{\eta}$, and $\tilde{\gamma}$ scaled proportionally from the AGILS parameters ${\alpha}$, ${\beta}$, ${\eta}$, and ${\gamma}$ using scales $\{1,1/2,1/3,1/4\}$. The optimal parameters that minimize the Error are selected as $\tilde{c}_{0}=1$, $p=0.2$, $\tilde{\gamma}=\gamma/3$, $\tilde{\alpha}=\alpha/4$, $\tilde{\beta}=\beta$, and $\tilde{\eta}=\eta$. MEHA terminates when $\mathrm{Error}<1/n$. For AGILS and its variants, the initial iterates are $x_{0}=[0,\dots,0]$, $y_{0}=\theta_{0}=a$, the parameters are set as $\epsilon=10^{-6}$, $c_{\tilde{y}}=50\sqrt{n}$ and $p_{0}=0.5,\ \varrho_{p}=0.02,\ c_{p}=1$. The step sizes $\alpha,\beta,\gamma$ are chosen as $\alpha_{k}=1/(L_{\psi_{x,k}}+c_{\alpha})$, $\beta_{k}=1/(L_{\psi_{y,k}}+c_{\beta})$, and $\gamma=1/(\rho_{f_{2}}+\rho_{g_{2}})$ where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ with the following values for the constants $L_{F_{x}}=L_{F_{y}}=L_{f_{x}}=\rho_{f_{1}}=\rho_{f_{2}}=L_{g_{1}}=0$, $\rho_{g_{1}}=\rho_{g_{2}}=1$, $L_{f_{y}}=n$, $c_{\alpha}=c_{\beta}=0.1$ and $\eta=1/(L_{f_{y}}+1/\gamma)$. AGILS and its two variants terminate when $\text{Error}<1/n$. For all experiments presented in this part, the approximations $\theta^{k}$ are computed at each iteration using the proximal gradient method to satisfy inexact criteria (17) and (20). For AGILS and its variants, the inexactness parameters are set as follows: $s_{k}=0.05/(k+1)^{1.05}$, $\tau_{k}=20/(k+1)^{0.7}$. All algorithms on the toy example are terminated if the runtime exceeds 200 seconds.
@@
Numerical results for two different problem dimensions, $n=200$ and $n=600$, are reported in Table 1. The results demonstrate that AGILS and its variants perform well in both computational efficiency and solution accuracy, consistently achieving the lowest Error with the shortest computational time. MEHA performs second, but it requires tuning of parameters from a set of candidate values, which can be a limitation. This limitation stems from MEHA’s stricter and less transparent step size rules, making tuning more important.
@@
Table 1: Comparisons of different methods on the toy example for dimensions $n=200$ and $n=600$. AGILS (Orig) refers to the original AGILS method. Its variants, AGILS_A and AGILS_R, are denoted as AGILS(Abs) and AGILS(Rel), respectively.
@@
Method Grid Random TPE MPCC IGJO VF-iDCA MEHA AGILS
Orig Abs Rel
Time(s) 0.59 5.89 29.61 20.99 6.06 0.52 0.09 0.07 0.08 0.08
Error 0.70 0.75 0.72 0.02 0.82 0.21 0.00 0.00 0.00 0.00